OpenStudy (anonymous):

What is the relationship between exponents and logarithms? How would you distinguish between the two, using both a graph and a sequence?

OpenStudy (amistre64):

log is a function that has the property: L(ab) = L(a) + L(b) exp is a function that has the property: E(a+b) = E(a) E(b) they are inverses of each other

OpenStudy (amistre64):

log(5*2) = log(5) + log(2) e^(5+2) = e^5 e^2 e^[log(5*2)] = e^[log(5) + log(2)] = e^[log(5)] e^[log(2)] = 5*2